Atlas Inc. is a toy bicycle manufacturing company producing a five inch small version of the bike that Lance Armstrong rode to win his first Tour de France. The assembly line at Atlas Inc. consists of seven work stations, each performing a single step. Stations and processing times are summarized here:

  • Step 1 (50 sec.): The plastic tube for the frame is cut to size.
  • Step 2 (40 sec.): The tube is put together.
  • Step 3 (55 sec.): The frame is glued together.
  • Step 4 (45 sec.): The frame is cleaned.
  • Step 5 (50 sec.): Paint is sprayed onto the frame.
  • Step 6 (65 sec.): Wheels are assembled.
  • Step 7 (60 sec.): All other parts are assembled to the frame.

Under the current process layout, workers are allocated to the stations as shown here:

  • Worker 1: Steps 1, 2
  • Worker 2: Steps 3, 4
  • Worker 3: Step 5
  • Worker 4: Step 6
  • Worker 5: Step 7

b. What is the capacity of this assembly line, in finished units/hour?

Capacity (units/hour)

c. What is the utilization of Worker 4, ignoring the production of the first and last units?

d. How long does it take to finish production of 80 units, starting with an empty process? (Round your final answer to two decimal places.)

Flow time (minutes)

g. Based on recommendations of a consultant, Atlas Inc. decides to reallocate the tasks among the workers to achieve maximum process capacity. Assume that if a worker is in charge of two tasks, then the tasks have to be adjacent to each other. Also, assume that the sequence of steps cannot be changed. What is the maximum possible capacity, in units per hour, that can be achieved by this reallocation? (Round your final answer to one decimal place.)

Max, Achievable Capacity (units/hour)
Answer #1

b.

Worker 2 is bottleneck and hence capacity = 3600/(55+45) = 36 units/hour

c.

Flow rate of worker 4 = 3600/65 = 55.38461538 units/hour

Utilization of worker 4 = 36/55.38461538 = 0.65

d.

1st unit will take = 50+40+55+45+50+65+60 = 365 seconds

and rest of the units will be processed with a cycle time of 100 seconds

So, time it takes to finish production of 80 units = 365+79*100 = 8265 seconds = 137.75 minutes

g.

Optimal production shedule :

Worker 1 : step 1
Worker 2 : step 2,3
Worker 3 : step 4,5
Worker 4 : step 6
Worker 5 : step 7

Now, worker 2 and 3 both are bottleneck with 55+40 = 95 seconds

So max achievable capacity = 3600/95 = 37.89473684 = 37.9 units/hour (Rounded to 1 decimal place)